Explain how the peaks in NMR spectrum correspond to the structure of isopentyl acetate.

Explain how the peaks in NMR spectrum correspond t
Answer #1
Concepts and reason

The concept used in this problem is based on Nuclear magnetic resonance (1HNMR)\left( {^{\rm{1}}{\rm{H NMR}}} \right) spectroscopy. The peaks in the NMR{\rm{NMR}} spectrum are used to identify the structure of the compound.

First, identify the number of signals of isopentyl acetate. After that assign the values of the chemical shift to identified signals.

Fundamentals

Principle of 1HNMR^{\bf{1}}{\bf{H NMR}} spectroscopy:

The principle of 1HNMR^{\rm{1}}{\rm{H NMR}} is that when the external magnetic field is applied, the energy is transferred from ground state to higher energy level and when the spin returns to ground state absorbed radiofrequency is emitted at the same frequency, that radiofrequency give the spectrum of the nucleus.

Chemical shift:

It is the difference between the resonance frequency of observed proton and tetramethylsilane (TMS)\left( {{\rm{TMS}}} \right) . TMS{\rm{TMS}} is used as a reference in NMR{\rm{NMR}} spectroscopy. The chemical shift of TMS{\rm{TMS}} is 0ppm{\rm{0 ppm}} .

Spin multiplicity:

The formula to calculate spin multiplicity is n+1n + 1 , here nn is a number of hydrogens of neighboring carbon.

Interpretation of the spectrum:

Firstly identify the different type of signals (protons).

Look at the multiplicity, it tells the number of protons of neighboring carbon.

Then see the position of the peaks, identify the chemical shift.

Some of the values of the chemical shift of 1HNMR^{\rm{1}}{\rm{H NMR}} are given as follows:

Type of proton
chemical
Approx.
shift (ppm)
0.9
Si(CHỊ),
-CH,
-CH-
-CH -
-C-C-CH, -
R-0-CH,
1.3
1.4
|
3.3-3.5
2.1-2.3
-CH

The structure of isopentyl acetate is as follows:

1H= multiplet
2H = triplet
6H = doublett
1 H2
CH
4
C.
3H = singlet
HzC? 3
5
.40
Сн, :
2H = quartet
isopentyl acetate

The peaks of 1HNMR^{\rm{1}}{\rm{H NMR}} corresponding to different protons of isopentyl acetate are as follows.

1H= 1.4 ppm
2H = 4.0 ppm
6H = 0.9 ppm
i H2
CH4C
3H = 2.0 ppm
Hac? 3
CH3
2H = 1.7 ppm
isopentyl acetate

The carbon (C5)\left( {{{\rm{C}}_{\rm{5}}}} \right) is directly attached to oxygen, so it is more deshielded. The value of the chemical shift of protons is upfield to 4.0ppm{\rm{4}}{\rm{.0 ppm}} . The carbon (C1)\left( {{{\rm{C}}_1}} \right) is directly attached to the carbonyl group, so it is also deshielded. The value of the chemical shift of protons is upfield to 2.0ppm{\rm{2}}{\rm{.0 ppm}} .

Ans:

The peaks of 1HNMR^{\bf{1}}{\bf{H NMR}} corresponding to different protons of isopentyl acetate are as follows:

1H= 1.4 ppm
2H = 4.0 ppm
6H = 0.9 ppm
i H2
CH4C
3H = 2.0 ppm
Hac? 3
CH3
2H = 1.7 ppm
isopentyl acetate